OpenStudy (anonymous):

Can someone give me one example about injective, surjective and biyective functions. I need to reinforce what I read.

OpenStudy (anonymous):

First of all I will start by stating what the synonyms for the terms are: Injective: one to one Surjective: onto

OpenStudy (anonymous):

A one to one function is such that The image of every element in the domain yields a distinct point in the range.

OpenStudy (anonymous):

Or formally: If \(f(x)\) is one to one then \[f(x_1) = f(x_2) \implies x_1 = x_2\]

OpenStudy (anonymous):

An onto function is simply one where the image of the domain is the entire target space.

OpenStudy (anonymous):

Finally bijective functions are functions that are both one to one and onto.

OpenStudy (anonymous):

Formally a function \(f:X->Y\) is onto if: For every \(y \in Y\) there is an \(x \in X\) such that \(f(x) = y\)

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Can you help me to determine if the following function is injective surjective, byyective or non of the above mencioned. f(x)=1/(x2 +1)

OpenStudy (anonymous):

no im helping my brother to get into the navy

OpenStudy (anonymous):

that is one of my brother questions

OpenStudy (anonymous):

Is it one to one? No, it is an even function so f(-x) = f(x) Therefore it is not injective Is it onto? Is the image of the domain the set of all real numbers? No it is not, so the function is not surjective.

OpenStudy (anonymous):

Just to confirm, this is the function \[f(x)=\frac{1}{x^2+1}\]correct?

OpenStudy (anonymous):

Correct me if im wrong, but if I want to find out if a function is 1-2-1 I should find out if this function is even? right?

OpenStudy (anonymous):

yeah, the function is correct

OpenStudy (anonymous):

No, even functions are not one to one, but not all functions that are not one to one are even.

OpenStudy (anonymous):

It is a sufficient condition, not a necessary one.

OpenStudy (anonymous):

ok till now I understand. can you give me another example like the one we just worked on, please

OpenStudy (anonymous):

Ok, for example, the function f(x)=\(x^3\) is both one to one and onto.

OpenStudy (anonymous):

can you show the work to determine that :)

OpenStudy (anonymous):

What is the image of the domain?

OpenStudy (anonymous):

The function \(x^3\) diverges to negative infinity for large \(-x\) and positive infinity for large \(x\).

OpenStudy (anonymous):

Then by the intermediate value theorem it will attain all the values between \(-\infty\) and \(\infty\). Therefore the image of the domain is \(\mathbb{R}\)

OpenStudy (anonymous):

ok got it. Tanks alchemista

OpenStudy (anonymous):

Also to prove its one to one, you need to show that if \(f(x_1)=f(x_2)\) then \(x_1 = x_2\)

OpenStudy (anonymous):

Show that if \((x_1)^3 = (x_2)^3\) then \(x_1 = x_2\). That should be easy.

OpenStudy (anonymous):

ok gonna try with other function later. Now I have to go to college. Thank you very much for you help. If I have any problem later Im gonna write on this threat . hope you dont mind?